HUBBLE SPACE TELESCOPE DAILY REPORT #1911 PERIOD COVERED: 0000Z (UTC) 07/01/97 - 0000Z (UTC) 07/02/97 Daily Status Report as of 183/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/1/2/3 7706 (NICMOS MULTIACCUM Dark Characterization 4/8) The purpose of this proposal is to obtain group 4 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. The observations completed with no reported problems. 1.3 Completed WF/PC-2 7450 (The Formation of Bulges of Spiral Galaxies: Central Visual-Infrared Color Gradients) The WF/PC-2 was used to make observations of the bulge in spiral galaxy NGC-5326 in order to obtain high-resolution optical and near infrared images for a complete survey sample of bulges of spiral galaxies. The observations were completed as scheduled, and no anomalies were reported. 1.4 Completed Two Sets of STIS/CCD 7700 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Telescope (CCD) was used to obtain slitless spectroscopy with the G750L grating to search for emission line galaxies at intermediate to high redshift. Field galaxy surveys have shown that at high redshift most galaxies have strong emission lines, and are compact and blue. This suggests that star formation rates were higher than at present,and that AGN may have been more frequent. The low background and high resolution of HST separates these objects from the background. The proposal completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrently with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.6 Completed Three Sets of NIC/1/2/3 7710 (NICMOS MULTIACCUM Dark Characterization 8/8) The purpose of this proposal is to obtain group 8 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. The observations completed with no reported problems. 1.7 Completed Four Sets of STIS/CCD 7312 (Investigating Volatile Depletion in Two Periodic Comets: 2P/Encke and 103P/Hartley 2) The Space Telescope Imaging Spectrograph (CCD) was used to make images of the periodic comet Enckke (2P) to investigate its volatile depletion. Current theories predict that the near-surface regions of cometary nuclei should become increasingly devolatilized during successive passages through the inner solar system. As reported in yesterday's report and in HSTAR 6109, there was a guide star acquisition failure at 182/0624Z, just before the second iteration of this proposal. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed Two Sets of NIC/1/2/3 7709 (NICMOS MULTIACCUM Dark Characterization 7/8) The purpose of this proposal is to obtain group 7 of the on-orbit darks for the MULTIACCUM predefined sequences requested by GOs, before the start of Cycle 7. As reported in yesterday's report and in HSTAR 6109, there was a guide star acquisition failure at 182/0901Z, just before the second iteration of this proposal. Otherwise, the observations completed with no reported problems. 1.9 Completed Two Sets of WF/PC-2 6905 (WF/PC-2 Cycle 6 Calibration: Internal Monitor) The WF/PC-2 was used to make several internal calibrations observations including biases and internal flats as part of a Cycle 6 calibrations procedure. A variety of internal exposures are run weekly to verify and monitor the health of the camera. As reported in yesterday's report and in HSTAR 6111, there was a guide star acquisition failure at 182/1035Z, just before the first iteration of this proposal. Otherwise, the observations were executed as scheduled, and no anomalies were noted. 1.10 Completed NIC/2 7327 (Elliptical Galaxy Evolution at High Redshift: IR Imaging of 60 Distant Clusters) The Near Infrared Camera (Camera Two) was used to make infrared observations high redshifted elliptical galaxies in galactic clusters CL1335+463 and CL1227+463. By measuring the structural properties of high redshift cluster elliptical galaxies, it is hoped to resolve the controversy regarding the age and formation history of bright cluster ellipticals. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 7063 (STIS SMOV CCD Contamination Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star GD153 in order to measure flux and look for any gross changes over time that might indicate contamination. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed Two Sets of NIC/1/2/3 7701 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used as part of a short NICMOS one orbit parallel program, to obtain "JHK" imaging of sky regions when another instrument is prime. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed WF/PC-2 7202 (The WF/PC-2 B Band Parallel Survey: A Systematic and Synoptic Study of Galaxy Formation and Evolution - Short Orbit Case) The WF/PC-2 was used to make observations as part of a B-Band Survey to make synoptic study of galaxy formation and evolution. These observations were taken in parallel with the prime science observations. They were completed as planned, and no anomalies were reported. 1.14 Completed Two Sets of NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-3277. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed FGS 6875 (Parallax of Delta Cephei) The Fine Guidance System was used to make astrometric observations of the Cephied class variable star Delta Cephei. Delta Cephei is a classical Cepheid variable, whose trignometric parallax is presently known to around 30%. HST observation shave the potential to reduce this uncertainty to 15%. The observations were completed as planned, and no anomalies were reported. 1.17 Completed NIC/1/2/3 7676 (NICMOS-One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to conduct an on-orbit calibration procedure. This was none in parallel with the prime science observations . The procedure was completed as scheduled, and no anomalies were noted. 1.18 Completed WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, while producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study using the WF/PC-2 of object OX169 will remedy this situation by undertaking redder, line-free observations of our three matched AGN samples. The proposal completed with no reported problems. 1.19 Completed STIS/CCD 7727 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to obtain parallel slitless spectroscopy with the G750L grating to search for emission line galaxies at intermediate to high redshift. Field galaxy surveys have shown that, at high redshift, most galaxies have strong emission lines, and are compact and blue. The observations were completed as planned, and no anomalies were reported. 1.20 Completed WF/PC-2 5946 (The Stellar Populations of Low Surface Brightness Galaxies) The WF/PC-2 was used to make observations of the low surface brightness galaxy UGC-12695. The objective of this program is to study the stellar population and evolution of this galaxy. The observations were completed as planned, and no anomalies were reported. 1.21 Completed Two Sets of NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make one orbit parallel calibration observations. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed STIS/CCD 7675 (STIS Non-Scripted Parallel Archive Proposal) The Space Telescope Imaging Spectrograph (CCD) was used in the parallel mode to obtain slitless spectroscopy with the G750L grating to search for emission line galaxies at intermediate to high redshift. The observations were completed as planned, and no anomalies were reported. 1.23 Completed STIS/CCD 7088 (STIS CCD Coronograph Occulting Bar Scattered Light) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the star HD60753 in order to study the scattered light profile from the occulting bars available for STIS imaging (50CORON) and spectroscopic apertures (52x0.2F1) using the CCD. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: Three guide star acquisition failures, reported in Section 2.1 of yesterday's report, are further described in 1.7, 1.8, and 1.9. As documented in HSTARs 6108 and 6110, there were two instances, at 182/064502Z and 182/083002Z, when the STOCC received notification that there had been a timed SPC busy reject error flag. Although there will be further analysis, initial indications are that the events are linked to guide star acquisiton failures occurring near the specified times. In each case, the STOCC send commands to clear the error flag. STIS EEPROM and EDAC memory dumps were performed at 182/1453Z. The STIS Bad Pixel Table was loaded at 182/1500Z. The SSR Interrupt Log was dumped at 183/0709Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH